# GeoBotsVR: A Robotics Learning Game for Beginners with Hands-on Learning Simulation

### 15 Feb 2024 v1

@@ 1. Introduction
In terms of mode of learning, researchers suggest that virtual reality (VR) enhances learning by increasing engagement and immersion (Hussein and Nätterdal, 2015; Dede, 2009; Mubarrat et al., 2020; Srinivasan et al., 2021). VR’s spatial navigation reduces the cognitive load in programming learning, surpassing traditional text-based methods (Lee and Wong, 2014). Additionally, VR provides a sense of self-presence (Ratan, 2013), fostering an embodied-cognitive learning experience (Melcer, 2018; Shin, 2017) that encourages more intuitive interaction with the content (Steed et al., 2016), potentially improving overall learning outcomes (Cheon and Grant, 2012). Therefore, I propose GeoBotsVR, an easily accessible and easy-to-play VR puzzle-solving/racing game designed to foster interest and motivation in robotics, electronics, and embedded systems of players through an enjoyable experience. In GeoBotsVR, the main gameplay focuses on players solving procedurally-generated puzzles using a customizable robot in different levels, while learning about robotics components, electronics, and programming along the way.
@@ 3.3. The Arcade: Main Gameplay
Players participate in the main gameplay using various components in The Arcade (Fig. 1b), which was designed by blending futuristic elements with 1980-90s arcade elements. Firstly, the player selects the game map from a globe menu simulating the Earth and nine different locations around the world (South American Mountains, African Desert, New York, East Coast Islands, European Snowlands, Russian Gulag, Asian Rainforest, Chinese Temple, and Japanese Industry) (Fig. 5a and b). After a map has been selected, the player uses the joysticks on the arcade machine to navigate their robot around the generated environment (Fig. 5c and d). Various other components of The Arcade assist the player in completing the level, such as the dock display providing in-game information (Fig. 5e), the top view display (Fig. 5f), the side view display (Fig. 5g), and various warning messages (Fig. 5h).

### 20 Feb 2024 v3

@@ 1. Introduction
In terms of mode of learning, researchers suggest that virtual reality (VR) enhances learning by increasing engagement and immersion (Hussein and Nätterdal, 2015; Dede, 2009; Mubarrat et al., 2020; Srinivasan et al., 2021; Chowdhury et al., 2022b, a). VR’s spatial navigation reduces the cognitive load in programming learning, surpassing traditional text-based methods (Lee and Wong, 2014). Additionally, VR provides a sense of self-presence (Ratan, 2013), fostering an embodied-cognitive learning experience (Melcer, 2018; Shin, 2017) that encourages more intuitive interaction with the content (Steed et al., 2016), potentially improving overall learning outcomes (Cheon and Grant, 2012). Therefore, I propose GeoBotsVR, an easily accessible and easy-to-play VR puzzle-solving/racing game designed to foster interest and motivation in robotics, electronics, and embedded systems of players through an enjoyable experience. In GeoBotsVR, the main gameplay focuses on players solving procedurally-generated puzzles using a customizable robot in different levels, while learning about robotics components, electronics, and programming along the way.
@@ 3.3. The Arcade: Main Gameplay
Players participate in the main gameplay using various components in The Arcade (Fig. 1b), which was designed by blending futuristic elements with 1980-90s arcade elements. Firstly, the player selects the game map from a globe menu simulating the Earth and nine different locations around the world (South American Mountains, African Desert, New York, East Coast Islands, European Snowlands, Russian Gulag, Asian Rainforest, Chinese Temple, and Japanese Industry) (Fig. 5a and b). After a map has been selected, the player uses the joysticks on the arcade machine to navigate their robot around the generated environment (Fig. 5c and d). There is a dock display which provides in-game information, such as score, time, robot stats, etc. (Fig. 5e). Various other components of The Arcade assist the player in completing the level, such as the top view display (providing a top-down view of the environment) (Fig. 5f), the side view display (providing an angled side view of the environment) (Fig. 5g), and various warning messages (Fig. 5h).